OpenStudy (anonymous):

Determine whether the function is even, odd, or neither h(x)=x^3-5

OpenStudy (lgbasallote):

hint: if ALL the exponents are odd then it is an odd function if ALL the exponents are even then it is an even function if NEITHER, then it's neither does that help?

jimthompson5910 (jim_thompson5910):

it also might help to think of h(x)=x^3-5 as h(x)=x^3-5x^0

OpenStudy (lgbasallote):

@jim_thompson5910 remind me...is 0 even or odd? it's neither even nor odd right?

jimthompson5910 (jim_thompson5910):

0 = 2*0

jimthompson5910 (jim_thompson5910):

so even

OpenStudy (anonymous):

its odd?

jimthompson5910 (jim_thompson5910):

what's odd 0? or the function?

OpenStudy (anonymous):

the function . .

jimthompson5910 (jim_thompson5910):

no, the exponent for the term -5, which is also -5x^0, is zero which is even

jimthompson5910 (jim_thompson5910):

so because of this clash, the function is neither even nor odd

OpenStudy (anonymous):

ohh , ok thanks (: